Circular No. 5877

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

X-RAY NOVA IN OPHIUCHUS Y. Tanaka, on behalf of the ASCA Team, communicates: “The hard x-

ray transient GRS 1716–249 = GRO J1719–24 was observed from ASCA on Oct. 5 UT. The source was detected near the edge of the SIGMA/GRANAT error circle (IAUC 5874). The position determined with ASCA is α = 17^{h}19^{m}35^{s}.2, δ = −25^{o}00′42′′ (equinox 2000.0), with an error-circle radius of 1′, coincident with that of the radio/optical candidate (cf. IAUC 5876) within 0′.5. The source intensity is ∼ 0.24 Crab in the range 0.5-10 keV, with rapid fluctuation of large amplitudes on time scales to < 1 s (similar to Cygnus X-1). The spectrum is well approximated by a single power law with a photon index of ∼ 1.6 and a hydrogen absorption column of ∼ 4 × 10^{21} atoms/cm^{2}.”

SUPERNOVA 1993ab IN NGC 1164 R. Kirshner and S. Gordon, Harvard-Smithsonian Center for Astro-

physics, report that a spectrogram (range 400-800 nm) of SN 1993ab obtained at the Multiple Mirror Telescope on Oct. 8 UT shows the object to be a typical type-Ia supernova at an age of ∼ 4 months. This is a different conclusion from that drawn on IAUC 5871.

GRO J1008–57 R. Petre and N. Gehrels, Goddard Space Flight Center, NASA, write:

“We have observed the newly discovered transient x-ray pulsar GRO J1008– 57 on Aug. 9 using the ROSAT PSPC. The source was detected at a count rate of 0.89 count/s in the band 0.1-2.4 keV. The best fit source position is α = 10^{h}09^{m}46^{s}, δ = −58^{o}17′32′′ (equinox 2000.0), with an error radius of 15′′, dominated by the systematics in the ROSAT attitude determination. The error circle is near the center of the BATSE error box (cf. IAUC 5838), and is consistent with the larger (1′ radius) ASCA error box (IAUC 5851). The best fit pulse period is 93.4 s. The source is ∼ 70 percent pulsed, with a well-defined double-peaked pulse profile. The 0.5- to 2.0-keV spectrum is most unusual; it is extremely soft: a fit to a power-law model requires an energy index in excess of 10, and a fit to a thermal bremsstrahlung model requires a temperature of 0.085 ± 0.02 keV. The column density is 3 × 10^{22} cm^{−2}, with a 90-percent confidence range of 2.4-4.7 × 10^{22} cm^{−2}. The observed 0.5- to 2.0-keV flux was 6.4 × 10^{−12} erg cm^{−2} s^{−1}.”

1993 October 9 Daniel W. E. Green